# Value-Penalized Auxiliary Control from Examples for Learning without Rewards or Demonstrations

### v1

@@ 3.2 Learning a Multitask Agent from Examples
where $y_{\mathcal{T}}$ are TD targets defined based on Eq. 2 with task-specific reward $\hat{R}_{\mathcal{T}}$.
Finally, a scheduler selects $\pi_{\mathcal{T}}$ to sample actions from during each environment interaction. Following [4, 31], the scheduler has a fixed period, allowing for a set number of policy switches in each episode. We use a weighted random scheduler and a small set of simple handcrafted high-level trajectories (e.g., reach then lift), following results from [4], where this approach performed as well as a more complex learned scheduler. The probability of choosing the main task is a hyperparameter $p_{\mathcal{T}_{\text{main}}}$, and each auxiliary task is chosen with probability $p_{\mathcal{T}_{k}}=(1-p_{\mathcal{T}_{\text{main}}})/K$. See Section D.6.2 for more scheduler details, including the full set of high-level trajectories.
3.3 Value Penalization in Example-Based Control
@@
A scheduled multitask agent will, by design, exhibit far more diverse behaviour than a single-task policy [4, 31]. We show in Section 4.2 that the buffer generated by this behavior, consisting of transitions resulting from multiple policies $\pi_{\mathcal{T}}$, can result in highly unstable and poorly calibrated Q estimates, especially in example-based control. In this section, we extend TD-error minimization, a fundamental component of many off-policy RL algorithms, with a penalty that encourages Q-function outputs to stay well-calibrated with respect to the reward model. Generally, the choice of the reward model and the loss function for estimating the Q-function can greatly impact learning efficiency (see Appendices A and B for details), but our value-penalty term applies to any reward model and commonly used regression-based loss functions. This penalty applies to both the single-task and multitask regime. For simplicity, we describe value penalization for the single-task framework.
Consider a reward model $\hat{R}(\cdot)$ where $\hat{R}(s^{*}),s^{*}\in\mathcal{B}^{*}$, indicates reward for successful states, while $\hat{R}(s),s\in\mathcal{B}$, for all other states. In AIL, $\hat{R}$ is a function of $D$, while in SQIL [6], for example, $\hat{R}(s^{*})=1$ and $\hat{R}(s)=0$. Assuming that $s^{*}$ transitions to itself, then for policy evaluation with the mean-squared error (MSE), we can write the TD target, $y:\mathcal{S}\times\mathcal{S}\to\mathbb{R}$, of Q-updates as $y(s,s^{\prime})=\hat{R}(s)+\gamma\mathbb{E}_{a^{\prime}}\left[Q(s^{\prime},a^{\prime})\right],$ (6) $y(s^{*},s^{*})=\hat{R}(s^{*})+\gamma\mathbb{E}_{a^{\prime}}\left[Q(s^{*},a^{\prime})\right],$ (7) where $(s,\cdot,s^{\prime})\sim\mathcal{B}$, $a^{\prime}\sim\pi(\cdot|s^{\prime})$, and $s^{*}\sim\mathcal{B}^{*}$. Eq. 7 can be replaced with $y(s^{*},s^{*})=\hat{R}(s^{*})$ if one considers successful states to be terminal, but this can cause bootstrapping errors when a task times out or does not terminate upon success [36], both of which are common practice in robotics environments. Regressing to TD targets Eqs. 6 and 7 will eventually satisfy the Bellman equation, but in the short term the targets do not satisfy $y(s,s^{\prime})\leq y(s^{*},s^{*})$. This is because the TD targets leverage bootstrapping of the current Q-estimate, an estimate that may not satisfy the Bellman equation and can exceed the bounds of valid Q-values, implying that approximation error of $Q$ updated with the MSE can be uncontrollable.
We introduce a simple resolution to this issue by adding a penalty to our TD updates for $s\in\mathcal{B}$ based on the current estimate of $\mathbb{E}_{s^{*}\sim\mathcal{B}^{*}}\left[V^{\pi}(s^{*})\right]$. This penalty term enforces the Q-estimate to focus on outputting valid values and drives the TD targets to satisfy the inequality $y(s,s^{\prime})\leq y(s^{*},s^{*})$. We introduce both a minimum and maximum value for $Q^{\pi}(s,a)$ respectively as
$Q^{\pi}_{\text{min}}=\hat{R}_{\text{min}}/(1-\gamma),$ (8) $Q^{\pi}_{\text{max}}=\mathbb{E}_{s^{*}\sim\mathcal{B}^{*}}\left[V^{\pi}(s^{*})\right],$ (9)
where $\hat{R}_{\text{min}}\leq\hat{R}(s)$ for all $s\in\mathcal{B}$. Then, the value penalty is
@@
$\mathcal{L}^{\pi}_{\text{pen}}(Q)=\lambda\mathbb{E}_{(s,a)\sim\mathcal{B}}\left[\left(\max(Q(s,a)-Q^{\pi}_{\text{max}},0)\right)^{2}+\left(\max(Q^{\pi}_{\text{min}}-Q(s,a),0)\right)^{2}\right],$ (10)
@@
where $\lambda\geq 0$ is a hyperparameter. Notice that when $\lambda\to\infty$, Eq. 10 becomes a hard constraint that is only satisfied by functions with valid outputs. It immediately follows that $y(s,s^{\prime})\leq y(s^{*},s^{*})$ holds with TD updates Eqs. 6 and 7.
We add value penalization $\mathcal{L}^{\pi}_{\text{pen}}(Q)$ to the MSE loss as a regularization term for learning the Q-function. In Appendix C, we show that the optimal solution of this regularized loss achieves low error on the unregularized loss. Consequently, if we use the regularized loss in policy evaluation, approximate policy iteration (API) still converges to the optimal solution. We state here the informal result that bounds the value error (see Theorem 1 in Appendix C for the formal theorem):
Informal Theorem.
Suppose the function class $\mathcal{Q}$ has bounded output and contains $Q^{\pi}$ for all $\pi$, and suppose we use the value-penalization loss for policy evaluation in API with $N$ samples for each iteration. Then, after $k$ iterations of API, we have that with probability at least $1-\delta$,
$\lVert Q^{*}-Q^{\pi_{k}}\rVert_{2,\rho_{0}}\leq\tilde{O}\left(\frac{1}{(1-\gamma)}\sqrt{\frac{\log(k/\delta)}{2N}}\right),$
where $\tilde{O}$ hides the constants and concentrability coefficients.
4 Experiments
Through our experiments, we seek to understand if VPACE improves stability and efficiency in example-based control. We also complete an ablation study of various hyperparameter options, and finally analyze the learned values for agents with and without value penalization.
4.1 Experimental Setup
(a) Samples of $\rho_{0}$, $\mathcal{B}^{*}_{\text{aux}}$, and $\mathcal{B}^{*}_{\text{main}}$.
(b) Real robot results.
Figure 3: Left: Examples from the initial state distribution $\rho_{0}$, auxiliary task example buffers $\mathcal{B}^{*}_{\text{aux}}$, and main task example buffers $\mathcal{B}^{*}_{\text{main}}$ for Unstack-Stack, Insert, sawyer_box_close, relocate-human-v0(-dp), RealDrawer, and RealDoor (see Section D.4 for all tasks). Right: Performance results for our two real robotic tasks.
@@ 4.1 Experimental Setup
We learn agents in a large variety of tasks and environments, including those originally used in LfGP [4] and RCE [3]. Specifically, the tasks from [4] involve a simulated Franka Emika Panda manipulator, a blue and green block, a fixed “bring” area for each block, and a small slot with $<$1 mm tolerance for inserting each block. The environment has a single, shared observation space and action space with multiple options for main and auxiliary tasks. We additionally study all tasks from [3]: a slightly modified subset of those from [11], involving a simulated Sawyer arm, and three of the Adroit hand tasks originally presented in [37]. We also generate three modified Adroit hand environments that use delta-position action spaces, instead of the absolute positions from the original environments, because we have found that policies learned in the original environments are very coarse and exploit simulator bugs (see Section D.2 for details). Finally, we study drawer and door opening tasks with a real Franka Emika Panda.
The reward models that we consider are a learned discriminator (Eq. 3, [5]), SQIL (Eq. 11 and Eq. 12, [6]), and RCE (Eq. 13 and Eq. 14, [3]). For each of these reward models, we can include value penalization (VP), auxiliary control from examples (ACE), or both (VPACE). VPACE-RCE and VP-RCE are not considered because RCE circumvents the issue of value penalization through its classification-based approach to updating value functions (see Section A.3).
All VPACE and ACE algorithms learn from examples of completed auxiliary tasks, in addition to the main task. In the simulated Panda environment, all main tasks use release, reach, grasp, and lift as auxiliary tasks, and each auxiliary task example dataset is reused between main tasks. In the Sawyer, Adroit, and real Panda environments, each main task uses reach and grasp as auxiliary tasks. Our scheduler approach for all environments is the same weighted random scheduler with a small set of handcrafted trajectories from [4]. All implementations are built on soft actor-critic (SAC) [20]. For more environment, algorithm, task, and implementation details, see Appendix D.
4.2 Main Task Performance Results
Figure 4: Top row: Performance results for four of the most challenging main tasks we considered. Bottom row: Average normalized performance for all main tasks we considered, separated by environment. A more detailed task performance breakdown can be found in Appendix E. Performance on auxiliary tasks is not shown. Methods that leverage ACE are shown with solid lines, and their single-task counterparts are shown with dashed lines. The shaded area corresponds to half standard deviation across five seeds (top), or a quarter standard deviation across tasks (bottom). To improve clarity, performance at each step is presented as an average across the five closest timesteps. Both value-penalization and auxiliary task exploration are beneficial for the most complicated tasks.
Our results for all algorithms for four of the most challenging main tasks we examined are shown at the top of Fig. 4, while the bottom shows normalized average results for all tasks, separated by environment. Our real Panda results are shown in Fig. 3(b). Since the Sawyer and Adroit tasks do not include specific success evaluation metrics, we only report returns. Policies are evaluated at 25k (environment step) intervals for 50 episodes for the simulated Panda tasks, 10k intervals for 30 episodes for the Sawyer and Adroit tasks, and 5k intervals for 10 episodes for the real Panda tasks.
Our results clearly show the benefits of combining auxiliary task exploration with value penalization. For most tasks, VPACE-SQIL learns faster and more stably than any other method. Notably, VPACE-SQIL can solve all tasks, apart from relocate-human-v0, by the final evaluation step. While VPACE-DAC also tends to perform quite well, it is less sample efficient than VPACE-SQIL on average, most likely owing to VPACE-SQIL’s reduced complexity and better discrimination between $\mathcal{B}$ and $\mathcal{B}^{*}$. For the Adroit non-DP tasks, VP-SQIL slightly outperforms methods including ACE, while SQIL alone performs comparably to VPACE methods, but we suspect that this is because these tasks have nuances allowing them to be “solved” with very coarse policies that may be undesirable for a real platform (see Section D.2 for details). ACE-RCE is consistently outperformed by both VPACE-SQIL and VPACE-DAC, and exhibits higher performance variance. We further examine this result, which conflicts with results presented in [3] showing very poor performance for both SQIL and DAC, in Appendix B. On our real tasks (RealDrawer and RealDoor), SQIL makes minor progress but is never able to solve either task, while VPACE-SQIL solves both.
Value penalization leads to a dramatic improvement in many of the more complex tasks, but presents comparable results to its exclusion in other tasks. Crucially, there are no tasks where its inclusion harms performance. VPACE-SQIL, VPACE-DAC, and ACE-RCE all outperform their single-task counterparts, particularly for the most complex tasks. The improvement for DAC in the simulated Panda environment is particularly stark; reflecting results from [4], DAC learns deceptive rewards for these tasks and, in turn, never learns to complete any of them except for Reach.
4.3 Ablations, Data Quantity, and Comparison to Full Trajectories and Sparse Rewards
(a) Ablation results.
(b) Q-value calibration with value penalization.
Figure 5: Left: Results for changes in data type, $\lambda$ value, and $\mathcal{B}^{*}_{\mathcal{T}}$ size. Right: Difference between Q-values and mean Q for example states for agents at 500k steps for a single episode rollout, without ACE on the left and with ACE on the right. Shaded regions shows half a standard deviation between seeds, while the dashed lines show the maximum output across the seeds. Value penalization ensures that the maximum stays below 0, verifying that $y(s,s^{\prime})\leq y(s^{*},s^{*})$.
We completed experiments with many variations from our original implementation of VPACE-SQIL in Unstack-Stack (see Fig. 5(a)). Our main experiments used $\lambda=10$ for value penalization strength, but we also tested $\lambda=\{1,100\}$, and found a negligible effect on performance, indicating robustness to $\lambda$ choice. We tested two different sizes of $\mathcal{B}^{*}_{\mathcal{T}}$ (for all tasks $\mathcal{T}$ in $\mathcal{T}_{\text{all}}$), $|\mathcal{B}^{*}_{\mathcal{T}}|=\{10,100\}$, while the main experiments used $|\mathcal{B}^{*}_{\mathcal{T}}|=200$ (following [3]). We found that $|\mathcal{B}^{*}_{\mathcal{T}}|=100$ had a negligible effect on performance, but $|\mathcal{B}^{*}_{\mathcal{T}}|=10$ slowed learning and impaired final performance. Even with $|\mathcal{B}^{*}_{\mathcal{T}}|=10$, VPACE-SQIL outperformed all baselines with $|\mathcal{B}^{*}_{\mathcal{T}}|=200$, apart from VPACE-DAC.
A natural question regarding our approach is how its performance compares to more traditional approaches, such as using full expert trajectories and inverse reinforcement learning (IRL), or using RL with true sparse rewards. To test the former, we added full trajectories to each $\mathcal{B}^{*}_{\mathcal{T}}$ (labelled +Full Trajectories in Fig. 5(a), +Full Trajectories & Actions for learning from actions as well, and VP-SQIL +Full Trajectories for single-task only), effectively making our approach similar to [4] but with value-penalization. Intriguingly, peak performance is reduced in this setting (especially without ACE), which we hypothesize is because the agent now has to minimize divergence between $\mathcal{B}$ and $\mathcal{B}^{*}_{\mathcal{T}}$ for many non-successful states, leading to an effect, commonly seen with dense reward functions, known as reward hacking [38]. This result suggests that inverse RL/adversarial IL can be significantly improved by switching to example-based control, but further investigation is required. To test RL with sparse rewards, we removed $\mathcal{B}^{*}_{\mathcal{T}}$ entirely, and instead use a ground truth $R(s,a)$ from the environment ($R(s,a)$ already having been used for evaluating success rate), similar to SAC-X [31] with value-penalization. SAC-X (Sparse Rewards) does not start accomplishing the main task at all in 500k environment steps, exhibiting substantially poorer performance than VPACE-SQIL and all other baselines (apart from DAC), indicating the high potential utility of immediately and consistently sampling expert success data.
4.4 Value Penalization and Calibration of Q-Values
@@
While our performance results show that value penalization improves performance, they do not explicitly show that $y(s,s^{\prime})\leq y(s^{*},s^{*})$ (see Section 3.3). To verify that this goal was met, we took snapshots of each learned agent at 500k steps and ran each policy for a single episode, recording per-timestep Q-values. Instead of showing Q-values directly, we show $Q(s_{t},a_{t})-\mathbb{E}_{s^{*}\sim\mathcal{B}^{*},a^{*}\sim\pi(\cdot|s^{*})}\left[Q(s^{*},a^{*})\right]$, which should be close to 00 when an episode is completed successfully, and should never climb above 00 for $y(s,s^{\prime})\leq y(s^{*},s^{*})$ to hold. We show the results of doing so in Unstack-Stack in Fig. 5(b) (see Section E.2 for other tasks).
Both VP-SQIL and VPACE-SQIL clearly do not violate $y(s,s^{\prime})\leq y(s^{*},s^{*})$, while both SQIL and ACE-SQIL do. ACE-SQIL, in particular, has no values, on average, where $y(s,s^{\prime})\leq y(s^{*},s^{*})$, indicating that it has learned poorly calibrated estimates for Q. This is reflected in our main performance results, where, in the most difficult Panda tasks in particular, the improvement from ACE-SQIL to VPACE-SQIL is pronounced. Consequently, scheduled auxiliary control can produce poorer policies than its single-task alternative unless it is coupled with value penalization.
5 Limitations
VPACE suffers from several limitations, though many of them are inherited from the use of reinforcement learning and learning from guided play. For an expansion of this section, including these inherited limitations, see Appendix F. In this work, we exclusively learn from numerical state data, rather than raw images, and raw images may be required for tasks involving objects that are not rigid. As well, we claim that example distributions are easier to generate than full expert trajectories, but for certain tasks, generating these example distributions may also be challenging. Finally, tasks we investigate in this work have roughly unimodal example success state distributions, and our method may not gracefully handle multimodality.
6 Conclusion

### v2

@@ III-B1 Intentions
where $y_{\mathcal{T}}$ are TD targets defined based on Eqs. 1 and 2 with task-specific reward $\hat{R}_{\mathcal{T}}$. Intuitively, each $\pi_{\mathcal{T}}$ aims to maximize the estimated task-specific value $Q_{\mathcal{T}}$.
III-B2 Scheduler
Figure 3: An example of fixed-period scheduler choices throughout an Unstack-Stack exploratory episode.
Given the $K+1$ intentions, a scheduler periodically selects a policy $\pi_{\mathcal{T}}$ to execute within an episode. Since each $\pi_{\mathcal{T}}$ aims to solve its own task, one can expect the transitions gathered composing by different $\pi_{\mathcal{T}}$ within the same episode to be diverse. In practice, we implement the scheduler to have a fixed period, guaranteeing a specific number of policy switches within each episode, depending on the episode horizon. We use a weighted random scheduler (WRS) with hyperparameter $p_{\mathcal{T}_{\text{main}}}$ where the probability of choosing the main task or an auxiliary task is $p_{\mathcal{T}_{\text{main}}}$ and $p_{\mathcal{T}_{k}}=(1-p_{\mathcal{T}_{\text{main}}})/K$, respectively. We combine a WRS with a small set of simple handcrafted high-level trajectories (e.g., reach then grasp then lift). The handcrafted trajectory definitions are reusable between main tasks, and are not required to use our framework. [6] demonstrated that this approach performed better than a more complex learned scheduler. At test time, the scheduler is unused, and only $\pi_{\text{main}}$ is evaluated.
III-C Value Penalization in Example-Based Control
@@
A scheduled multitask agent exhibits far more diverse behaviour than a single-task agent [5, 6]. We show in Figs. 8 and 9 that the buffer generated by this behavior, consisting of transitions resulting from multiple policies, can result in highly overestimated Q-values in EBC. This overestimation leads the policy to maximize an incorrect objective. In this section, we propose a novel penalty for TD algorithms that encourages Q-estimates to stay within the range of valid returns with respect to the reward model. This penalty applies to both the single-task and multitask regime. For simplicity, we describe value penalization for the former.
Figure 4: Samples from initial-state distribution $\rho_{0}$, auxiliary task examples $\mathcal{B}^{*}_{\text{aux}}$, and main task examples $\mathcal{B}^{*}_{\text{main}}$ for all tasks. The simulated Panda tasks additionally share $\mathcal{B}^{*}_{\text{release}}$ and $\mathcal{B}^{*}_{\text{lift}}$, while the Adroit original and dp environments share data since $\mathcal{S}$ does not change.
Notice that regressing to TD targets Eqs. 1 and 2 will eventually satisfy the Bellman equation, but in the short term the TD targets do not satisfy $y(s,s^{\prime})\leq y(s^{*},s^{*})$. This is because the TD targets are computed by bootstrapping from a Q-estimate that may not satisfy the Bellman equation and can exceed the bounds of valid Q-values, implying that approximation error of $Q$ updated via MSE can be uncontrollable. We resolve this issue with a penalty for our TD updates for $s\in\mathcal{B}$. We add a minimum and a maximum $Q^{\pi}(s,a)$ as $Q^{\pi}_{\text{min}}=\hat{R}_{\text{min}}/(1-\gamma)$ and $Q^{\pi}_{\text{max}}=\mathbb{E}_{\mathcal{B}^{*}}\left[V^{\pi}(s^{*})\right]$, where $\hat{R}_{\text{min}}\leq\hat{R}(s)$ for all $s\in\mathcal{B}$. Then, the value penalty is defined to be
@@
$\begin{split}\mathcal{L}^{\pi}_{\text{pen}}(Q)=\lambda\mathbb{E}_{\mathcal{B}}[&\left(\max(Q(s,a)-Q^{\pi}_{\text{max}},0)\right)^{2}\\ &+\left(\max(Q^{\pi}_{\text{min}}-Q(s,a),0)\right)^{2}],\end{split}$ (5)
@@
where $\lambda\geq 0$ is a hyperparameter. When $\lambda\to\infty$, Eq. 13 becomes a hard constraint. It immediately follows that $y(s,s^{\prime})\leq y(s^{*},s^{*})$ holds with TD updates Eqs. 1 and 2. We add value penalization $\mathcal{L}^{\pi}_{\text{pen}}(Q)$ to the MSE loss as a regularization term for learning the Q-function.
IV Experiments
We aim to answer the following questions through our experiments: (RQ1) How does the sample efficiency of VPACE compare to EBC, inverse RL, and exploration-bonus baselines? (RQ2) How important is it to include value penalization (VP) with ACE? (RQ3) How does our value penalty compare to existing Q regularizers? (RQ4) How does VPACE compare to algorithms that use full trajectories and true sparse rewards?
IV-A Experimental Setup
IV-A1 Environments
@@
We conduct experiments in a large variety of tasks and environments, including those originally used in LfGP [6] and RCE [3]. Specifically, the tasks in [6] involve a simulated Franka Emika Panda arm, a blue and green block, a fixed “bring” area for each block, and a small slot with $<$1 mm tolerance for inserting each block. This environment provides various manipulation tasks that share the same state-action space. The tasks in [3] are a modified subset of those from [11], involving a simulated Sawyer arm, and three of the Adroit hand tasks originally presented in [33]. We also generate three modified delta-position (dp) Adroit hand environments, because we found that policies learned in the original absolute-position environments lack finesse and exploit simulator bugs. Finally, we study drawer and door opening tasks with a real Franka Emika Panda.
IV-A2 Baselines
Figure 5: Sample efficiency performance plots for main task only for all main tasks. Performance is an interquartile mean (IQM) across 5 timesteps and 5 seeds with shaded regions showing 95% stratified bootstrap confidence intervals [7]. Other baselines (RCE, DAC, SQIL+RND) are shown only in Fig. 6 for clarity. The use of ACE significantly improves upon SQIL, and the addition of VP (VPACE) resolves instability issues that occur in Stack, Unstack-Stack, Bring, Insert, and hammer-human-v0-dp.
Figure 6: Average performance across all simulated main tasks, separated by environment, including baselines not shown in Fig. 5.
We consider many baselines, including an approach based on recursively classifying examples (RCE, [3]), a learned reward model for off-policy RL (DAC, [14]), a defined reward model for off-policy RL (SQIL, [15]), and a combination of the best performing baseline (SQIL) with a method that provides an exploration bonus to unseen data (SQIL+RND, [34, 35]). Finding that SQIL significantly outperformed the baselines, we use it as the default reward model for our main experiments, although VP and ACE can be added to any of the approaches described above. Specifically, in our results, VPACE refers to SQIL with both VP and ACE applied, while ACE refers to SQIL with ACE and without VP applied.
IV-A3 Implementation
Observation space, action space, and auxiliary task details are shown in Table II. All simulated tasks use 200 examples per task, while the real world tasks use 50 examples per task. All implementations are built on LfGP [6, 36] and SAC [19]. For more environment and implementation details, see our open-source code.
TABLE II: Environment details.
Obs. Space Act Space Aux. Tasks
Sim. Panda [6, 26] pos, vel, grip pos, prev. grip pos, obj. pos, obj. vel XYZ delta-pos, binary grip reach, grasp, lift, release
Sawyer [3, 11] pos, obj. pos, grip pos, 3 frame stack XYZ delta-pos, binary grip reach, grasp
Adroit [3, 33] pos, vel, obj. pos, finger pos (up to) 6-DOF abs. / delta-pos, finger pos reach, grasp
Real Panda pos, obj. pos (ArUco [37]), grip pos, 2 frame stack X / XY delta-pos, binary grip reach, grasp
IV-B Performance Results
IV-B1 Main Task VP and ACE Benefits
To answer RQ1, we compare VPACE with existing EBC, inverse RL, and exploration-bonus approaches on all tasks and evaluate their success rates. Since the Sawyer and Adroit tasks do not evaluate success, we only report returns. Policies are evaluated at 25k (environment step) intervals for 50 episodes for the simulated Panda tasks, 10k intervals for 30 episodes for the Sawyer and Adroit tasks, and 5k intervals for 10 episodes for the real Panda tasks. Fig. 5 shows that VPACE has significant improvement over other approaches in both sample efficiency and final performance, particularly for the most difficult tasks, such as Unstack-Stack and Insert. We can observe that VPACE can consistently solve all tasks, while other EBC methods have significantly wider confidence intervals, especially in the Panda environment. Another notable observation is that SQIL+RND is unable to improve upon SQIL. We speculate that the exploration bonus term actually diverts the policy from solving the main task in order to maximize the intrinsic return.
IV-B2 ACE without VP
We also include ACE in Fig. 5 to address RQ2. In the simulated Panda environment, ACE performance significantly degrades as training continues. As elaborated on in Section IV-C1, we hypothesize that this is correlated to value overestimation being exacerbated by executing multiple policies. For Sawyer and Adroit environments, VP does not provide benefit (i.e., VPACE and ACE have similar performance), which might be due to the tasks being simpler and having smaller observation spaces than the Panda tasks. Nonetheless, VP does not harm performance and including auxiliary tasks provides a clear benefit.
IV-B3 Real Robot Performance
The sample-efficiency improvement gained via VPACE allowed us to test it on real-life tasks (two right columns of Fig. 4), where the bottom right plots in Fig. 5 demonstrate that VPACE can solve real-world tasks in only a few hours of real execution time, whereas SQIL does not solve either task. Our real robot tasks execute at 5 Hz, and, with a small amount of time to allow for environment resetting, 1000 environment steps corresponds to roughly 4 real minutes, meaning RealDrawer is learned by VPACE in about 100 minutes, while RealDoor is learned in about 200 minutes. The real world examples of success, for both main tasks and auxiliary tasks, are collected in less than a minute. Our open-source code contains further implementation details for our real world setup.
IV-C Analysis and Additional Results
IV-C1 Q-Value Overestimation and Value Penalization
Figure 7: Results for different example-based approaches to regularizing Q estimates. We measure performance using return instead of success rate as the alternatives achieved no success at all.
@@ IV-C1 Q-Value Overestimation and Value Penalization
To verify that VP enforces $y(s,s^{\prime})\leq y(s^{*},s^{*})$, we took snapshots of each learned Unstack-Stack and Insert agent, for both VPACE and ACE, at 300k steps and ran each policy for a single episode, recording per-timestep Q-values. Instead of showing Q-values directly, in Figs. 8 and 9, we show $Q(s_{t},a_{t})-\mathbb{E}_{s^{*}\sim\mathcal{B}^{*}}\left[V(s^{*})\right]$, which should be at most 00 for $y(s,s^{\prime})\leq y(s^{*},s^{*})$ to hold. ACE clearly violates $y(s,s^{\prime})\leq y(s^{*},s^{*})$, while VPACE does not. Furthermore, Fig. 9 shows examples of the consequences of overestimated Q-values: for $(s_{t},a_{t})$ pairs that appear to be out-of-distribution (OOD), $Q(s_{t},a_{t})-\mathbb{E}_{s^{*}\sim\mathcal{B}^{*}}\left[V(s^{*})\right]>0$, and the resulting policy reaches these states instead of the true goal.
To understand the importance of the $y(s,s^{\prime})\leq y(s^{*},s^{*})$ constraint, we investigate other choices of regularization techniques (RQ3), in particular a L2 regularizer on Q-values (C2F, [31]) and a regularizer that penalizes the Q-values of out-of-distribution actions (CQL, [21]). Fig. 7 indicates that using other existing regularization techniques does not enable ACE to perform well. We suspect that the L2 regularizer may be too harsh of a penalty on all learning, while [21] does not have any penalization on the magnitudes of Q-values in general, meaning that they can potentially still have uncontrollable bootstrapping error.
IV-C2 Comparison to Full Demonstrations and True Rewards
A practitioner may ask how our method compares to using traditional forms of feedback, such as using full-expert trajectories and inverse reinforcement learning (IRL), or using RL with true sparse rewards (RQ4). To test the former, we include two variants: a buffer containing both 200 examples and 200 $(s,a)$ pairs from expert trajectories of states only (+Trajs), and the same for both states and actions (+Trajs & Acts), resulting in two variants of LfGP [6] with VP. For sparse rewards, we use the provided sparse-reward function from the task (SAC-X), which is scheduled auxiliary control [5] with VP, although VP is based on the maximum theoretical return in these experiments, rather than $y(s,s^{\prime})\leq y(s^{*},s^{*})$ since $s^{*}$ are not available to SAC-X.
Figure 8: Difference between Q-values and expected value for example states for partially trained agents (snapshot from 300k environment steps) for a single episode rollout. The dashed lines show the maximum output across the seeds, and the shaded regions show the standard deviation between seeds. VP ensures that the maximum stays below 0, verifying that $y(s,s^{\prime})\leq y(s^{*},s^{*})$.
Figure 9: $Q(s_{t},a_{t})-\mathbb{E}_{\mathcal{B}^{*}}\left[V^{\pi}(s^{*})\right]$ for parts of episodes used to generate the results from Fig. 8. It appears that the use of ACE without VP results in overestimated Q-values for out-of-distribution (OOD) $(s_{t},a_{t})$ pairs (e.g., grasping both blocks simultaneously, or nearly inserting the incorrect block), resulting in highly suboptimal policies. VPACE, on the other hand, correctly learns to label a potentially OOD state with low value.
Figure 10: Results for changes in the form of feedback. SAC-X indicates the use of true environment sparse rewards only.
Figure 11: Variations of success example quantity and $\lambda$ value.
Fig. 10 shows that the peak performance is reduced when learning with expert trajectories. We hypothesize that the divergence minimization objective leads to an effect, commonly seen with dense reward functions, known as reward hacking [38]. This effect was also previously shown to occur in inverse RL [6]. In short, the agent receives the same positive reward for reaching intermediate states in the demonstration as it does for reaching final states, despite the fact that intermediate states may not lead to the optimal final states. This result suggests that inverse RL/adversarial IL can be significantly improved by switching to example-based control. Learning with a sparse-reward function only does not accomplish the main task at all in either Unstack-Stack nor Insert, exhibiting substantially poorer performance than all other baselines. This result matches previous research in which it has been shown that a demonstration buffer can significantly improve sample efficiency in RL [25, 39]. Our results show that an example buffer may provide a similar effect, and may even be a more optimal form of guidance than full demonstrations.
IV-C3 Expert Data Quantity and $\lambda$ Value
Finally, to examine the robustness of our approach, we also compare various values for the quantity of examples of success (200 / task for main experiments) as well as the value of $\lambda$, which controls the strength of value penalization (10.0 for main experiments). Fig. 11 shows that changing the value of $\lambda$ to 1 and 100 has no effect on performance. Dropping from 200 to 100 examples has only a minor negative impact on performance in Insert, while the drop to 10 examples has a significant negative effect. However, even with only 10 examples, performance on Unstack-Stack still far exceeds all other baselines.
V Conclusion